OpenStudy (anonymous):

use one sided limits at x->3 from the right and left to fin the lim x^2-9/(abs(x-3))

OpenStudy (anonymous):

I will do the limit from the left side, and leave to you the other part: \[\lim_{x \rightarrow 3^-}{x^2-9 \over \left| x-3 \right|}=\lim_{x \rightarrow 3^-}{(x+3)(x-3) \over -(x-3)}=\lim_{x \rightarrow 3^-}-(x+3)=-6\]

OpenStudy (anonymous):

ok so in the denom what do you do for the right side?

OpenStudy (anonymous):

like how did you know to put a - sign in front?

myininaya (myininaya):

if we are looking to left of 3, then we are looking at numbers smaller than 3 x<3=>|x-3|=-(x-3)

myininaya (myininaya):

if you are looking to the right of 3, then we are looking at numbers greater than 3 x>3=>|x-3|=x-3

OpenStudy (anonymous):

oh alright i have never done it that way